OpenStudy (anonymous):

2log_4(10)− log_4(5^2) hey guys is the answer 1 or 1.43...

OpenStudy (anonymous):

@ParthKohli @Bookworm14

jhonyy9 (jhonyy9):

2log_4(10) = log_4(10^2) = log_4(100) = log_4(4*5^2) = log_4(4) + log_4(5^2) log_4(4)=1 1 + log_4(5^2) -log_4(5^2)= 1 so the 1 will be right sure